OpenStudy (31356):

Can you check this? (Factoring using the Distributive Property) MEDAL REWARDED!!

OpenStudy (rebeccapink6398):

could you please come to my question and help?

OpenStudy (31356):

t^2 h+3t=t(ht+3)

OpenStudy (31356):

Sure

OpenStudy (31356):

81r+48rs=3r(27+16s)

OpenStudy (anonymous):

First distribute the t to (ht+3) t(ht) = ? t(3) = ?

OpenStudy (31356):

82e^3-122ef=2e(41e^2-61f)

OpenStudy (31356):

@Mertsj @jim_thompson5910

OpenStudy (anonymous):

Lets do one at a time...

OpenStudy (31356):

Please check if it's correct.

jimthompson5910 (jim_thompson5910):

They are all correct 31356, nice work

OpenStudy (31356):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome